# Unusual intra-burst variations of polarization states in FRB 20210912A and FRB 20230708A : Effects of plasma birefringence?

### v1

@@ 5.3 Effects of Linear Birefringence
Propagation through a linear birefringent medium (LBM), which introduces a phase difference between its two orthogonal linear polarization modes, may also induce significant circular polarization in (originally) linearly polarized emission. We assume that initially the emission is fully linearly polarized with a position angle of $\gamma(t)$ – i.e. $\gamma$ is the angle between the electric field on the sky plane and an arbitrarily chosen reference axis. The intrinsic time dependence of $\gamma$ can be attributed to the properties of the emission source (see Section 6.1). An LBM introduces a phase difference $\delta$ between the two orthogonal linear components of the electric field along its optical axes (i.e. the ‘fast’ axis and the ‘slow’ axis). The angle between the reference axis and the fast axis of the LBM is denoted by $\phi$. Here we consider a scenario where the LBM is close to the FRB emission site, while the interstellar medium (ISM) and the intergalactic medium (IGM) only introduce Faraday rotation of the PA of the linear polarization by a frequency dependent angle, $\alpha(\lambda)$. The observed polarization vector, expressed in terms of $\chi$ (PA) and $\psi$ (EA), follows the relation (see Appendix B)
@@ 6 Great Circles from Linear Birefringence
Figure 12: Variation of the inferred intrinsic PA ($\gamma=\Gamma-\phi$) across sub-burst $A$ [A, left panel] and sub-burst $B$ [B, right panel] of FRB 20210912A. Corrections for $\rm RM=4.55\;rad\>m^{2}$ have been applied. The thick dashed lines represent the inferred RVM with $\alpha=80.2^{\circ}$, $\Theta_{A}=62.5^{\circ}$ and $\beta_{A}=17.7^{\circ}$. The dotted line in the right panel shows the dashed line shift by $14^{\circ}$. See Section 6.1 for details.
@@ 6.1 Motion along the Great Circle
Figure 11 shows the motion along the great circle — i.e. the temporal evolution of $\Gamma$ — across the primary sub-bursts of FRB 20230708A and FRB 20210912A in different frequency sub-bands. In FRB 20230708A, the first peak does not show any frequency dependence of $\Gamma$, although the second peak exhibits measurable differences among the four sub-bands. The change in $\Gamma$ at the second peak — with respect to its frequency independent value at the maximum of the first peak — is larger at lower frequencies. We note that the first peak of FRB 20230708A is almost fully linearly polarized and exhibits an roughly constant RM, while the second peak has a varying RM as well as significant circular polarization. The primary sub-burst of FRB 20210912A exhibits frequency-dependent motion along the great circle, varying RM as well as significant circular polarization – the same features as the second peak of FRB 20230708A. Presence of circular polarization, apparent variation of RM and frequency dependence of $\Gamma$ thus appear to be connected, and have the same underlying reason. The observed frequency dependence of $\Gamma$ is not naturally explained by the LBM scenario. Under the assumption that $\phi$ is frequency independent, the observed frequency dependence of $\Gamma$ may be attributed to the frequency dependence of the intrinsic PA ($\gamma$) — possibly due to propagation effects through a different medium before encountering the LBM (e.g. the magnetosphere of a compact source), or different frequencies being emitted from different regions and thus tracing differently orientated local magnetic fields.
@@ 7.2 Possible Linear Birefringent Media
in the strong magnetic field regime (cyclotron frequency much larger than the frequency of the radiation), where $e$, $m_{e}$ and $n_{e}$ are the charge and the mass and the number density of electrons and $\Delta l$ is the thickness of the plasma shell. We note that, in this scenario, the magnetic field associated with this plasma has to be predominantly perpendicular to the line of sight for the natural polarization modes to be approximately linear. Such an orientation of the magnetic field naturally arises in the outer magnetosphere or the near-wind region of a neutron star (Lyutikov, 2022). In case of FRB 20210912A, if the two sub-bursts are associated with emission from different regions of a compact object (e.g. a neutron star), different orientations of the local magnetic field (and hence the axes of the LBM) in the outer magnetosphere (or the near-wind region) would naturally explain the observed difference in the Poincaré sphere trajectories for the two sub-bursts.
Assuming a pair plasma in a high magnetic field such that the cyclotron frequency is much higher than the frequency of the electromagnetic waves, we find that a plasma screen with a column density of $\sim 10^{10}-10^{11}$ cm-2 may lead to the observed inclinations ($\Delta_{0}$) of the great circle trajectories on the Poincaré sphere. This column density is not unreasonable for a relatively thin plasma screen near the light cylinder of a neutron star (e.g. Beloborodov & Thompson, 2007; Thompson et al., 2002). As the phase difference between the two natural polarization modes varies linearly with the wavelength of the electromagnetic radiation in this scenario, this is also consistent with the lack of significant frequency dependence of $\Delta_{0}$ within the observing frequency band.
@@ 8 Summary
We investigated temporal variation of the polarization states of two fast radio bursts, FRB 20230708A and FRB 20210912A, which show time-dependent conversion between linear and circular polarizations. The polarization vector was found to trace great circle trajectories on the surface of the Poincaré sphere for both FRBs, with the two sub-bursts of FRB 20210912A tracing different great circles. The trajectories of the polarization vector show weak frequency dependence within the observing frequency band. The observed temporal variation of the polarization vector for these two FRBs are unlikely to have originated from instrumental effects. These polarization features may be signatures of a transition between two partially coherent orthogonal polarization modes or propagation through a birefringent medium. However, we find that transition between two orthogonal polarization modes or effects of propagation through an elliptically birefringent medium are unlikely to give rise to the observed polarization behaviour. The observed variations of the polarization states of these two FRBs are qualitatively consistent a magnetospheric origin of the bursts and the effects of propagation through a birefringent medium with linearly polarized modes in the outer magnetosphere or near-wind region of a neutron star.

### v2

@@ 5.3 Effects of Linear Birefringence
Propagation through a linear birefringent medium (LBM), which introduces a phase difference between its two linearly polarized orthogonal propagation modes, may also induce significant circular polarization in (originally) linearly polarized emission. We assume that the incident radiation is fully linearly polarized with a position angle of $\gamma(t)$ – i.e. $\gamma$ is the angle between the electric field on the sky plane and an arbitrarily chosen reference axis. The intrinsic time dependence of $\gamma$ can be attributed to the properties of the emission source (see Section 6.1). An LBM introduces a phase difference $\delta$ between the two orthogonal linear components of the electric field along its optical axes (i.e. the ‘fast’ axis and the ‘slow’ axis). The angle between the reference axis and the fast axis of the LBM is denoted by $\phi$. Here we consider a scenario where the LBM is close to the FRB emission site, while the interstellar medium (ISM) and the intergalactic medium (IGM) only introduce Faraday rotation of the PA of the linear polarization by a frequency dependent angle, $\alpha(\lambda)$. The observed polarization vector, expressed in terms of $\chi$ (PA) and $\psi$ (EA), follows the relation (see Appendix B)
@@ 6 Great Circles from Linear Birefringence
Figure 12: Variation of the inferred intrinsic PA ($\gamma=\Gamma-\phi$) across sub-burst $A$ [A, left panel] and sub-burst $B$ [B, right panel] of FRB 20210912A. Corrections for $\rm RM=4.55\;rad\>m^{2}$ have been applied. The thick dashed lines represent the inferred RVM with $\alpha=80.2^{\circ}$, $\Theta_{A}=62.5^{\circ}$ and $\beta_{A}=17.7^{\circ}$. The dotted line in the right panel shows the dashed line shifted by $14^{\circ}$. See Section 6.1 for details.
@@ 6.1 Motion along the Great Circle
Figure 11 shows the motion along the great circle — i.e. the temporal evolution of $\Gamma$ — across the primary sub-bursts of FRB 20230708A and FRB 20210912A in different frequency sub-bands. In FRB 20230708A, the first peak does not show any frequency dependence of $\Gamma$, although the second peak exhibits measurable differences among the four sub-bands. The change in $\Gamma$ at the second peak — with respect to its frequency independent value at the maximum of the first peak — is larger at lower frequencies. We note that the first peak of FRB 20230708A is almost fully linearly polarized and exhibits a roughly constant RM, while the second peak has a varying RM as well as significant circular polarization. The primary sub-burst of FRB 20210912A exhibits frequency-dependent motion along the great circle, varying RM as well as significant circular polarization – the same features as the second peak of FRB 20230708A. Presence of circular polarization, apparent variation of RM and frequency dependence of $\Gamma$ thus appear to be connected, and have the same underlying reason. The observed frequency dependence of $\Gamma$ is not naturally explained by the LBM scenario. Under the assumption that $\phi$ is frequency independent, the observed frequency dependence of $\Gamma$ may be attributed to the frequency dependence of the intrinsic PA ($\gamma$) — possibly due to propagation effects through a different medium before encountering the LBM (e.g. the magnetosphere of a compact source), or different frequencies being emitted from different regions and thus tracing differently orientated local magnetic fields.
@@ 7.2 Possible Linear Birefringent Media
in the strong magnetic field regime (cyclotron frequency much larger than the frequency of the radiation), where $e$, $m_{e}$ and $n_{e}$ are the charge and the mass and the number density of electrons and $\Delta l$ is the thickness of the plasma shell. We note that, in this scenario, the magnetic field associated with this plasma has to be predominantly perpendicular to the line of sight for the natural polarization modes to be approximately linear. Such an orientation of the magnetic field naturally arises in the outer magnetosphere or the near-wind region of a neutron star (Lyutikov, 2022). In case of FRB 20210912A, if the two sub-bursts are associated with emission from different regions of a compact object (e.g. a neutron star), different orientations of the local magnetic field (and hence the axes of the LBM) in the outer magnetosphere (or the near-wind region) would naturally explain the observed difference in the Poincaré sphere trajectories for the two sub-bursts. This scenario is consistent with a magnetospheric origin of FRBs, supporting the conclusion of some recent studies (e.g. Niu et al., 2024; Nimmo et al., 2025).
Assuming a pair plasma in a high magnetic field such that the cyclotron frequency is much higher than the frequency of the EM waves, we find that a plasma screen with a column density of $\sim 10^{10}-10^{11}$ cm-2 may lead to the observed inclinations ($\Delta_{0}$) of the great circle trajectories on the Poincaré sphere. This column density is not unreasonable for a relatively thin plasma screen near the light cylinder of a neutron star (e.g. Beloborodov & Thompson, 2007; Thompson et al., 2002). As the phase difference between the two natural polarization modes varies linearly with the wavelength of the EM radiation in this scenario, this is also consistent with the lack of significant frequency dependence of $\Delta_{0}$ within the observing frequency band.
As mentioned previously, no further bursts have been reported from either FRB 20210912A or FRB 20230708A to date. However, detection of more bursts from these sources in future could potentially provide an excellent opportunity to test the hypothesis described in this section and to probe the magnetic field properties of the emitting objects.
@@ 8 Summary
We investigated temporal variation of the polarization states of two fast radio bursts, FRB 20230708A and FRB 20210912A, which show time-dependent conversion between linear and circular polarizations. The polarization vector was found to trace great circle trajectories on the surface of the Poincaré sphere for both FRBs, with the two sub-bursts of FRB 20210912A tracing different great circles. The trajectories of the polarization vector show weak frequency dependence within the observing frequency band. The observed temporal variation of the polarization vector for these two FRBs are unlikely to have originated from instrumental effects. These polarization features may be signatures of a transition between two partially coherent orthogonal polarization modes or propagation through a birefringent medium. However, we find that transition between two orthogonal polarization modes or effects of propagation through an elliptically birefringent medium are unlikely to give rise to the observed polarization behaviour. The observed variations of the polarization states of these two FRBs are qualitatively consistent with a magnetospheric origin of the bursts and the effects of propagation through a birefringent medium with linearly polarized modes in the outer magnetosphere or near-wind region of a neutron star.